OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2?

OpenStudy (anonymous):

use again the formula (y-y1) = m(x-x1) like what you have did in the previous one :)

OpenStudy (anonymous):

K

OpenStudy (anonymous):

is it y+8=2(x-6)?

OpenStudy (anonymous):

it should be -2 :)

OpenStudy (anonymous):

for the value of your slope

OpenStudy (anonymous):

oh so y+8=-2(x-6)?

OpenStudy (anonymous):

yep and then expand it more :)

OpenStudy (anonymous):

K thanks